# A predictive model for bubble–particle collisions in turbulence

### v1

@@ 5.2.2 Collision rate predictions and implications
While the collision kernel facilitates comparison between scenarios with different bubble and particle number densities, it is not the most direct comparison in practice when $r_{b}$ varies since it implies that the gas holdup $\varepsilon_{g}\propto r_{b}^{3}$, which may not be the case (Yianatos et al., 2010). This is an important detail, especially for the flotation of fine particles. In this context, the mineral recovery rate is low due to reduced collision efficiency at small $r_{p}$, so smaller bubbles are employed to mitigate this problem (Ahmed and Jameson, 1985; Miettinen et al., 2010). The collision kernel then takes a lower value as has already been observed in figure 8(a). However, considering only $\Gamma$ in this scenario ignores the positive effect on the overall collision rate due to the increase in the bubble number density assuming a constant gas volume flow rate. Therefore, in figure 9 we plot $\Gamma n_{b}$ (i.e. the collision rate per particle) for constant $\varepsilon_{g}$ at a typical value of 10% (Yianatos et al., 2010). As expected, the shape of the curves are identical to those in figure 8(a); though crucially, $\Gamma n_{b}$ increases when $r_{b}$ decreases. This shows that using smaller bubbles has a net positive effect on the collision rate, and our model prediction supports the usage of smaller bubbles for flotation, which is consistent with results in the literature (Ahmed and Jameson, 1985; Miettinen et al., 2010).
@@ 6 Conclusion
In this study, we develop a model that predicts a priori the bubble–particle collision rate of inertial particles in homogeneous isotropic turbulence given the bubble, particle, and turbulence properties: namely, the bubble radius, the bubble density, the particle radius, the particle response time, the liquid kinematic viscosity, the liquid density, the Taylor Reynolds number, the dissipation rate, and gravity. This is achieved by adopting the frozen turbulence approximation proposed by Jiang and Krug (2025), such that the existing parametrisations of the bubble–particle collision efficiency in still fluid can be applied. The main input required for this approach is the bubble slip speed p.d.f.. We assume independent bubble slip velocity components and model their probability density functions using normal distributions with the mean of the vertical component given by Ruth et al. (2021) and Liu et al. (2024) (i.e. (16)), and the standard deviation given by Berk and Coletti (2024) (i.e. (28)). The resulting bubble slip speed p.d.f.s are compared with finite-size bubble simulations over $1\leq 1/Fr\leq 10$ with $0.5\leq St_{b}\leq 6.3$ and $Re_{\lambda}=64$. Good agreement is observed for $1/Fr\geq 4$. For $St_{b}=2.8$ and $6.3$, we additionally calculate the bubble–particle collision rates from the simulations and find satisfactory agreement with the model predictions. This result is consistent with our proposed requirement that $T_{L}^{w}/\tau_{f}\gtrsim 1$ for the frozen turbulence approximation, which suggests that the model should be valid for $1/Fr\geq 4$ when $0.5\leq St_{b}\leq 6.3$ and $Re_{\lambda}=64$. Applying our model to typical flotation parameters confirms that the collision rate can be raised by using larger particles, smaller bubbles, or stronger turbulence. Importantly, we find that even for fine particles, particle inertia can play a significant role in determining the overall collision rate.
@@
In summary, our model includes the particle inertia effect on the collision rate and provides significantly better predictions of the bubble–particle collision rate in homogeneous isotropic turbulence than existing theories, even at the $St_{p}\to 0$ limit. Considering the high inhomogeneity of the turbulence in actual flotation cells, our model can be used on a subgrid-scale level in industrial computer fluid dynamics simulations to predict the local collision rate, at least for smaller bubble and particle Stokes numbers where the behaviour of the dispersed phases is determined by the local flow field. Potential future improvements of the model include a more accurate representation of $E_{c}^{(in)}(St_{p})$, to more faithfully capture the dependence of the collision rate on $St_{p}$; as well as incorporating the turbulence-induced spatial segregation of bubbles and particles (Calzavarini et al., 2008a; Chan et al., 2023, 2025), which can reduce the collision rate especially when $St_{p}\approx 1$ (Jiang and Krug, 2025).
@@ Appendix A Fitting parameters for the inertial collision efficiency
In this appendix, we provide the expressions used to evaluate $\Gamma^{(Kos)}$ for completeness. Rewriting (1.2),
@@ B.1 Definition of $N_{1,2,3}$
$N_{1}=3YFU_{T},\quad N_{2}=2XFU_{T}+u_{p}\mathfrak{f},\quad N_{3}=u_{pt}-YFU_{T}.$ (40)
Here, $u_{p}=2r_{p}^{2}(\rho_{p}/\rho_{f}-1)\mathfrak{g}/(9\nu)$ is the particle terminal velocity assuming Stokes drag ($u_{p}=0$ in §4 since the simulated particles do not experience gravity), $u_{pt}=(r_{p}/\tau_{\eta})/(30\pi)^{1/2}$ is the collision velocity attributed to local fluid shear,
@@
$\mathfrak{f}$ =$=\begin{cases}\,\alpha/2&\text{for $\alpha<1$}\\ \,0.5+0.5[1-\exp(-0.85(\alpha-1))]&\text{for $\alpha\geq 5$.}\end{cases}$ (45)
@@ Appendix D Grid convergence test
To ensure that the collision kernel is not sensitive to the grid spacing, we perform an additional simulation for the highest $Re_{b}$ case ($(St_{b},1/Fr)=(6.3,10)$) at twice the grid resolution. Figure 12 shows that the ratio of the collision kernels obtained is almost unity, which suggests that the grid resolution is sufficient.
Figure 12: (a) The collision kernels at the standard grid resolution $\Gamma$ and that at the doubled grid resolution $\Gamma^{\prime}$, and (b) the ratio of Γ/$\Gamma/\Gamma^{\prime}$ as a function of the particle Stokes number for $(St_{b},1/Fr)=(6.3,10)$.

### v2

@@ 5.2.3 Collision rate predictions and implications
While the collision kernel facilitates comparison between scenarios with different bubble and particle number densities, it is not the most direct comparison in practice when $r_{b}$ varies since it implies that the gas holdup $\phi\propto r_{b}^{3}$, which may not be the case (Yianatos et al., 2010). This is an important detail, especially for the flotation of fine particles. In this context, the mineral recovery rate is low due to reduced collision efficiency at small $r_{p}$, so smaller bubbles are employed to mitigate this problem (Ahmed and Jameson, 1985; Miettinen et al., 2010). The collision kernel then takes a lower value as has already been observed in figure 9(a). However, considering only $\Gamma$ in this scenario ignores the positive effect on the overall collision rate due to the increase in the bubble number density assuming a constant gas volume flow rate. Therefore, in figure 10 we plot $\Gamma n_{b}$ (i.e. the collision rate per particle) for constant $\varepsilon_{g}$ at a typical value of 10% (Yianatos et al., 2010). As expected, the shape of the curves are identical to those in figure 9(a); though crucially, $\Gamma n_{b}$ increases when $r_{b}$ decreases. This shows that using smaller bubbles has a net positive effect on the collision rate, and our model prediction supports the usage of smaller bubbles for flotation, which is consistent with results in the literature (Ahmed and Jameson, 1985; Miettinen et al., 2010).
@@ 6 Conclusion
In this study, we develop a model that predicts a priori the bubble–particle collision rate of inertial particles in homogeneous isotropic turbulence given the bubble, particle, and turbulence properties: namely, the bubble radius, the bubble density, the particle radius, the particle response time, the liquid kinematic viscosity, the liquid density, the Taylor Reynolds number, the dissipation rate, and gravity. Our model accounts for bubble behaviour in turbulence as well as the distorted flow field around finite-size bubbles. A summary of our model is displayed in figure 11. As figure 11 shows, our model adopts the frozen turbulence approximation proposed by Jiang and Krug (2025), such that the existing parametrisations of the bubble–particle collision efficiency in still fluid can be applied. The main input required for this approach is the bubble slip speed p.d.f.. We assume independent bubble slip velocity components and model their probability density functions using normal distributions with the mean of the vertical component given by Ruth et al. (2021) and Liu et al. (2024) (i.e. (14)), and the standard deviation given by Berk and Coletti (2024) (i.e. (26)). The resulting bubble slip speed p.d.f.s are compared with finite-size bubble simulations over $1\leq 1/Fr\leq 10$ with $0.5\leq St_{b}\leq 6.3$ and $Re_{\lambda}=64$. Good agreement is observed for $1/Fr\geq 4$. For $St_{b}=2.8$ and $6.3$, we additionally calculate the bubble–particle collision rates from the simulations and find satisfactory agreement with the model predictions. This result is consistent with our proposed requirement that $T_{L}^{w}/\tau_{f}\gtrsim 1$ for the frozen turbulence approximation, which suggests that the model should be valid for $1/Fr\geq 4$ when $0.5\leq St_{b}\leq 6.3$ and $Re_{\lambda}=64$. Applying our model to typical flotation parameters confirms that the collision rate can be raised by using larger particles, smaller bubbles, or stronger turbulence. Importantly, we find that even for fine particles, particle inertia can play a significant role in determining the overall collision rate.
Figure 11: Summary of the proposed model. The numbers in brackets refer to the expression of the corresponding quantities.
@@
In summary, our model includes the particle inertia and particle settling effects on the collision rate and provides significantly better predictions of the bubble–particle collision rate in homogeneous isotropic turbulence than existing theories, even at the $St_{p}\to 0$ limit. Considering the high inhomogeneity of the turbulence in actual flotation cells, our model can be used on a subgrid-scale level in industrial computer fluid dynamics simulations to predict the local collision rate, at least for smaller bubble and particle Stokes numbers where the behaviour of the dispersed phases is determined by the local flow field. Presently, the expressions for the collision efficiency and the mean vertical slip velocity presented in (9) and (14) are for fully contaminated bubbles. In principle though, our general modelling approach (i.e. the frozen turbulence approach and the modelling of the bubble slip speed p.d.f. with independent normally distributed components) is not restricted to such bubbles, as long as the corresponding expressions of the collision efficiency and the mean vertical slip velocity are used. Potential future improvements of the model include a more accurate representation of $E_{c}^{(in)}(St_{p})$, to more faithfully capture the dependence of the collision rate on $St_{p}$; as well as incorporating the turbulence-induced spatial segregation of bubbles and particles (Calzavarini et al., 2008a; Chan et al., 2023, 2025), which can reduce the collision rate especially when $St_{p}\approx 1$ (Jiang and Krug, 2025). In addition, interactions with neighbouring bubbles may influence the collision kernel in practice, but these effects are not currently represented in the model and should be addressed in future research.
@@ Appendix A Fitting parameters for the inertial collision efficiency
In this appendix, we provide the expressions used to evaluate $\Gamma^{(Kos)}$ for completeness. This model is based on the combined surface and velocity integral
$\Gamma^{(Kos)}$ $=$ $\oiint_{S_{c}}\iiint\limits_{-\infty}^{\infty}\int_{0}^{\infty}\Lambda(-\Delta v_{r}\cdot f(w_{bo,x})f(w_{bo,y})f(w_{bo,z})f(v_{pt})$
$\mathrm{d}v_{pt}\mathrm{d}w_{bo,x}\mathrm{d}w_{bo,y}\mathrm{d}w_{bo,z})\mathrm{d}S,$
[…]
$f(\cdot)$ is the probability density function (p.d.f.), $w_{bo,x,y,z}$ are components of the bubble slip velocity with respect to the fluid velocity in the absence of the bubble, and $v_{pt}$ is the relative bubble–particle velocity due to local velocity gradients. Rewriting (B),
@@ B.1 Definition of $N_{1,2,3}$
$N_{1}=3YFU_{T},\quad N_{2}=2XFU_{T}-v_{s}\mathfrak{f},\quad N_{3}=v_{pt}-YFU_{T}.$ (40)
Here, $v_{s}=2r_{p}^{2}(1-\rho_{p}/\rho_{f})\mathfrak{g}/(9\nu(1+0.169Re_{p}^{2/3}))$ is the particle terminal velocity with non-Stokesian drag correction, $v_{pt}=(r_{p}/\tau_{\eta})/(30\pi)^{1/2}$ is the collision velocity attributed to local fluid shear,
@@
$\mathfrak{f}$ $=\begin{cases}\,\alpha/2&\text{for $\alpha<1$}\\ \,0.5+0.5[1-\exp(-0.85(\alpha-1))]&\text{for $\alpha\geq 1$.}\end{cases}$ (45)
@@ Appendix D Grid convergence test
To ensure that the collision kernel is not sensitive to the grid spacing, we perform an additional simulation for the highest $Re_{b}$ case ($(St_{b},1/Fr)=(6.3,10)$) at twice the grid resolution for non-settling particles. Figure 14 shows that the ratio of the collision kernels obtained is almost unity, which suggests that the grid resolution is sufficient.
Figure 14: (a) The collision kernels at the standard grid resolution $\Gamma$ and that at the doubled grid resolution $\Gamma^{\prime}$, and (b) the ratio of $\Gamma/\Gamma^{\prime}$ as a function of the particle Stokes number for $(St_{b},1/Fr)=(6.3,10)$.